# Crop Row Switching for Vision-Based Navigation: A Comprehensive Approach for Efficient Crop Field Navigation

### v1

@@ III Methodology
Figure 1 portrays a state machine representing the process of crop row switching, spanning across a total of seven states. Table I describe each state of the process given in Figure 1 encountered during row switching manoeuvre. The robot was driven within the crop rows using Triangle Scan Method (TSM) based crop row navigation framework [6] we have developed in our previous work up to state $A$. The switching manoeuvre is composed of three steps: row exit, U-turn and re-entry. The transitions $A\rightarrow B\rightarrow C$ belong to the step of row exit. The U-turn step contains to the transitions $C\rightarrow D\rightarrow E\rightarrow F$ while the transition $F\rightarrow G$ is considered as re-entry step. The methods and techniques used in each of these three steps are explained in sections III-B, III-C and III-D respectively.
TABLE I: States encountered during crop row switching manoeuvre.
@@
$A$
Position of the robot during initial EOR detection.
$B$
Robot reaches the EOR position.
$C$
Robot enters the headland area completely passing EOR position.
$D$
Robot turns 90°towards the direction of next crop row from state $C$.
$E$
Robot aligns itself with the next crop row traversing forward from state $D$.
$F$
Robot turns 90°facing itself towards the starting point of the next crop row while residing within the headland buffer.
$G$
Robot enters the next crop row.
III-A Robotic Platform
A Mark-1 robot from Hexman Robotics was used to test the proposed crop row switching algorithm. Figure 2 shows the robotic setup with a front mounted Intel RealSense D435i camera and a Reach RS+ RTK GNSS receiver. A NVIDIA Jetson AGX Orin developer kit was used as the onboard computer for the robot. Mark-1 is a skid steer robot with a ground clearance of 128 mm. The dimensions of the robot are $526\times 507\times 244$ mm.
Figure 2: Hexman Mark-1 robot in the Sugar Beet Field.
III-B Row Exit Step
Row exit is the process in which the robot drives itself completely out of the currently traversing crop row after detecting the EOR. The EOR is initially detected at state $A$ where the robot estimates the relative 3D coordinate of the starting point of the next crop row robot would enter. The Y value of this relative 3D coordinate would represent the inter-row distance $d_{r}$ between current and the next crop row. The $d_{r}$ value is useful when robot is traversing during the $D\rightarrow E$ transition. The robot transits through the sates $A\rightarrow B\rightarrow C$ after the re-entry point detection using a combination of visual feature matching and odometry as explained in following subsections.
III-B1 Re-entry Point Detection
The re-entry point detection module is extended based on the TSM crop row detection pipeline from our previous work [5] on crop row detection. Similar to TSM, a deep learning based skeleton segmentation of crop rows was used as the input for the re-entry point detector. As illustrated in Figure 3, the ROI $AL_{2}L_{3}B$ was defined if the next intended turn is to the left ($AR_{2}R_{3}C$ for right). Points $A,B$ and $C$ were determined using the ”Anchor scans” step of TSM [5]. The horizontal green line in Figure 3 was obtained using the EOR detector [6]. Equation 1 yields a point $P_{t}$ by scanning for the pixel sum along $AP$ line where point $P$ is an arbitrary point on $L_{2}L_{3}B$ path. Similarly, equation 2 yields a point $A_{t}$ by scanning the pixel sum along $\overline{A}P_{t}$ line where point $\overline{A}$ is an arbitrary point on $AL_{1}$ line. The intersection between the EOR line and $A_{t}P_{t}$ is identified as the re-entry point $R$ for the next crop row. The depth information from the corresponding depth image was used to determine the 3D coordinate corresponding to the point $R$ which was then used to determine $d_{r}$.
@@ III-B1 Re-entry Point Detection
III-B2 A to B Transition
The TSM based crop row navigation framework from our previous work [6] was proven to be able to maintain the average heading of the robot relative to the crop row under 1°. Therefore, the relative heading angle between the robot and the crop row was assumed to be under 1°at state $A$. The detected EOR line in Figure 3 demarcates the headland area and the crop row region within the RGB image obtained from the front mounted camera. The RGB image was cropped below the EOR line and saved as a reference image $I_{R}$ at state $A$. The robot was then moved towards the EOR with a constant forward linear velocity while calculating the local feature (Scale Invariant Feature Transform [23] was used) similarity score in each new image captured by the robot camera with $I_{R}$. The robot was stopped and assumed to reach state $B$ when the feature similarity score dropped below a experimentally determined threshold. The threshold value was determined by observing the feature similarity score while driving the robot to the actual EOR position using teleoperation.
III-B3 B to C Transition
@@ III-C U-Turn Step
The U-turn step involves the robot taking two 90°turns with a linear navigation stage ($D\rightarrow E$) in between. ”Headland buffer” region is the space in the headland that is directly in front of a given crop row bound by the EOR, edge of the field and the two centre lines of the inter-row space between adjacent crop rows. The goal of the U-turn step is to bring the robot into the headland buffer region of the next crop row while facing towards the next crop row to be traversed. It was experimentally verified that the TSM can resume it’s normal crop row navigation framework from this point(state $F$) onward to traverse in to the crop row. The 90°turns and $D\rightarrow E$ transition is executed with wheel odometry feedback of the robot. The practical behaviour of the robot during rotation stages ($C\rightarrow D$ and $E\rightarrow F$) ensured that the robot would reach the headland buffer when the $D\rightarrow E$ transition distance was set to $d_{r}$.
III-D Re-Entry Step
@@
The robot is within the headland buffer region at state $F$ and state $G$ is reached when the robot enters the next crop row to be traversed. The re-entry step is the transition from state $F$ to state $G$ where the robot moves from the headland buffer into the crop row in front of it. This transition was realized by launching the TSM based crop row following framework at state $F$. The TSM was able to detect the crop row in front of the robot and navigate the robot into the crop row. This behaviour of TSM based re-entry navigation is verified by the experiment outlined in Section IV-C.
IV Results and Discussion
@@
The row switching manoeuvre presented in this paper is a three-step process that involves 6 state transitions as illustrated in Figure 1. The accuracy and the performance of each transition is individually evaluated during the experiments. The Hexman Mark-1 robot was programmed to follow a crop row and automatically detect the EOR and re-entry positions. The proposed row-switching manoeuvre is automatically started based on the EOR detection trigger. The path of the robot during the row-switching manoeuvre was tracked using the onboard RTK GNSS tracker with sub-centimeter accuracy.
@@
An experiment was set up in a selected area of 10 crop rows within a real sugar beet field. The GNSS coordinates of the 10 crop rows were recorded with sub-centimetre accuracy by driving the robot through each crop row at very slow speeds by an expert human driver. These ground truth GNSS coordinates of each crop row was then used to generate a regression line which would be used as a reference to calculate errors in autonomous navigation during row switching manoeuvre of the robot. The robot was allowed to autonomously execute the row switching manoeuvre among the 10 crop rows turning in both directions (left and right turns). A total of 18 row-switching trials were conducted during this experiment as plotted in Figure 5. The GNSS coordinates were converted to Universal Transverse Mercator (UTM) coordinates for plotting and error calculations. The errors during each state transition of each row switching manoeuvre trial are illustrated in Figure 4.a where distance errors and angular errors are normalized within a scatter plot. There are 9 x-ticks in Figure 4.a between each pair of adjacent states which represents a pair of consecutive real-world crop rows where the trial took place. A box and whisker plot of the errors for each transition is denoted in Figure 4.b. This normalised representation provides a comparative illustration of the error magnitudes in each state transition. The plots corresponding to translational steps stands above the zero line while box and whisker plots of angular transitions are centered close to zero. Table II presents the median errors of traversal during each stage of the row-switching manoeuvre. Normalized median percentage error ($\alpha=E_{median}/E_{max,T}$) was calculated where $E_{max,T}$ is the maximum absolute error for the type $T$ (T being a distance or an angle) across the entire manoeuvre and $E_{median}$ is the median error for each transition. The vision based transition $A\rightarrow B$ records the highest $\alpha$ value representing the transition with highest error. The remaining transitions record relatively lower $\alpha$ values representing accurate navigation relative to the vision-based transition.
@@
TABLE II: Linear and angular errors during each transition of the row switching maneuver.
@@
IV-A Crop row exit and headland entry
The row exit step ($A\rightarrow B\rightarrow C$ ) in the proposed row switching manoeuvre could be identified with the densely distributed GNSS coordinates in Figure 5 leading towards headland from within the crop row in each trial. The dense distribution of GNSS coordinates attributes to the slower speed of row switching manoeuvre relative to the in-row navigation speed of the crop row navigation algorithm.
The row crop row exit ($A\rightarrow B$ transition) is a vision-based navigation stage where the headland entry ($B\rightarrow C$ transition) uses the wheel odometry to guide the robot into the headland area. This difference in feedback modalities is reflected in the distance errors of each transition. The visual feedback in $A\rightarrow B$ transition ensures that the robot has successfully reached the EOR position with visual confirmation. Therefore, it is vital for successfully reaching the EOR despite the higher error margins compared to wheel odometry. Majority of the errors in $A\rightarrow B$ and $B\rightarrow C$ transitions are positive, which indicates that the robot always travel further into the headland area beyond desired positions at states $B$ and $C$. This trend doesn’t have significantly adverse impact on the overall row switching manoeuvre since such extra distance traversed into the headland wouldn’t cause the robot to damage crops during the U-turn step.
The robot would stop at 52.6 cm ($L_{robot}$) away from the actual EOR position in an ideal $C$ state. However, the overall maximum error in the row exit step ($E_{ABC,max}$) was recorded as 64.27 cm, a distance which the robot would move further away from the desired position at state C. Based on these observations, the minimum width $W_{H,min}$ for the headland space was calculated to be 143.17 cm using Equation 3. The coefficient of $L_{robot}$ term in Equation 3 was set to $1.85$ since the RTK-GNSS receiver used to measure the robot motion was mounted at 45 cm behind the front of the robot. This coefficient must be changed to $(1+R)$ where $R$ depends on the position of the RTK-GNSS receiver on the robot, if this experiment is being repeated on a different robot. The robot was expected to be aligned with the crop row at state $A$ within a heading error margin of 2°. A heading error beyond 2°at state $A$ leads the robot to cross into the headland buffer of an adjacent crop row at state $C$, which would cause it to skip 1 crop row during switching or re-enter to the same crop row it traversed as seen in the unsuccessful attemps in Figure 5.

### v2

@@ 3.5 Crop Row Switching
Figure 6 portrays a state machine representing the process of crop row switching, spanning across a total of seven states. Table 1 describes each state of the process given in Figure 6 encountered during the row switching manoeuvre. The blue colour overlay in Figure 3 illustrates the crop row switching manoeuvre executed by the robot upon detection of the EOR (state $A$). The robot uses the EOR detector described in [27] to detect the EOR while traversing the crop row. The robot is switched back to row-following mode at the end of the row-switching manoeuvre (state $G$). The switching manoeuvre is composed of three steps: row exit, U-turn and re-entry. The transitions $A\rightarrow B\rightarrow C$ belong to the step of row exit. The U-turn step contains the transitions $C\rightarrow D\rightarrow E\rightarrow F$ while the transition $F\rightarrow G$ is considered as a re-entry step. The methods and techniques used in each of these three steps are explained in sections 3.5.1, 3.5.2 and 3.5.3 respectively.
The experiment on row switching manoeuvre outlined in Section 4.2 reported a lower success rate due to the slipping of the robot within the headland area during the 90 turns ($C\rightarrow D$ and $E\rightarrow F$) of the row switching manoeuvre. The row-switching algorithm is improved by adding translational odometry tracking to the state transitions $C\rightarrow D$ and $E\rightarrow F$ to accurately traverse the inter-row distance $D\rightarrow E$. The impact of this improvement is reported in the experimental results reported in Section 4.3.
Figure 6: Crop row switching state machine. $d_{r}$: Distance between current and next crop row, A: Initial Detection of the EOR, B: Robot is at the EOR, C: Robot traversed a distance equal to its length into the headland, D: Robot turn 90 towards the next row direction from state C, E: Robot traverse $d_{r}$ distance forward from state D, F: Robot turn 90 towards the next row direction from state E, G: Robot re-enter the next crop row.
Table 1: States encountered during crop row switching manoeuvre.
@@
$A$ Position of the robot during initial EOR detection.
$B$ Robot reaches the EOR position.
$C$ Robot enters the headland area completely passing EOR position.
$D$ Robot turns 90 towards the direction of the next crop row from state $C$.
$E$ Robot aligns itself with the next crop row traversing forward from state $D$.
$F$ Robot turns 90 facing itself towards the starting point of the next crop row while residing within the headland buffer.
$G$ Robot enters the next crop row.
3.5.1 Row Exit Step
Row exit is the process in which the robot drives itself completely out of the currently traversing crop row after detecting the EOR. The EOR is initially detected at state $A$ where the robot estimates the relative 3D coordinate of the starting point of the next crop row the robot would enter. The Y value of this relative 3D coordinate would represent the inter-row distance $d_{r}$ between the current and the next crop row. The $d_{r}$ value is useful when the robot is traversing during the $D\rightarrow E$ transition. The robot transits through the states $A\rightarrow B\rightarrow C$ after the re-entry point detection using a combination of visual feature matching and odometry as explained in the following subsections.
Re-entry Point Detection:
The re-entry point detection module is extended based on the TSM crop row detection pipeline from our previous work [28] on crop row detection. Similar to TSM, a deep learning-based skeleton segmentation of crop rows was used as the input for the re-entry point detector. As illustrated in Figure 7, the ROI $AL_{2}L_{3}B$ was defined if the next intended turn is to the left ($AR_{2}R_{3}C$ for right). Points $A,B$ and $C$ were determined using the "Anchor scans" step of TSM [28]. The horizontal green line in Figure 7 was obtained using the EOR detector [27]. Equation 2 yields a point $P_{t}$ by scanning for the pixel sum along $AP$ line where point $P$ is an arbitrary point on $L_{2}L_{3}B$ path. Similarly, equation 3 yields a point $A_{t}$ by scanning the pixel sum along $\overline{A}P_{t}$ line where point $\overline{A}$ is an arbitrary point on $AL_{1}$ line. The intersection between the EOR line and $A_{t}P_{t}$ is identified as the re-entry point $R$ for the next crop row. The depth information from the corresponding depth image was used to determine the 3D coordinate corresponding to the point $R$ which was then used to determine $d_{r}$.
@@ 3.5.1 Row Exit Step
A to B Transition:
The TSM-based crop row navigation framework was proven to be able to maintain the average heading of the robot relative to the crop row under 1 [27] . Therefore, the relative heading angle between the robot and the crop row was assumed to be under 1 at state $A$. The detected EOR line in Figure 7 demarcates the headland area and the crop row region within the RGB image obtained from the front-mounted camera. The RGB image was cropped below the EOR line and saved as a reference image $I_{R}$ at state $A$. The robot was then moved towards the EOR with a constant forward linear velocity while calculating the local feature (Scale Invariant Feature Transform [18] was used) similarity score in each new image captured by the robot camera with $I_{R}$. The robot was stopped and assumed to reach state $B$ when the feature similarity score dropped below an experimentally determined threshold. The threshold value was determined by observing the feature similarity score while driving the robot to the actual EOR position using teleoperation.
B to C Transition:
@@ 3.5.2 U-Turn Step
The U-turn step involves the robot taking two 90 turns with a linear navigation stage ($D\rightarrow E$) in between. "Headland buffer" region is the space in the headland that is directly in front of a given crop row bound by the EOR, edge of the field and the two centre lines of the inter-row space between adjacent crop rows. The goal of the U-turn step is to bring the robot into the headland buffer region of the next crop row while facing towards the next crop row to be traversed. It was experimentally verified that the TSM can resume its normal crop row navigation framework from this point(state $F$) onward to traverse in to the crop row. The 90 turns and $D\rightarrow E$ transition is executed with wheel odometry feedback of the robot. The practical behaviour of the robot during rotation stages ($C\rightarrow D$ and $E\rightarrow F$) ensured that the robot would reach the headland buffer when the $D\rightarrow E$ transition distance was set to $d_{r}$.
3.5.3 Re-Entry Step
@@
The robot is within the headland buffer region at state $F$ and state $G$ is reached when the robot enters the next crop row to be traversed. The re-entry step is the transition from state $F$ to state $G$ where the robot moves from the headland buffer into the crop row in front of it. This transition was realized by launching the TSM-based crop row following framework at state $F$. The TSM was able to detect the crop row in front of the robot and navigate the robot into the crop row. This behaviour of TSM-based re-entry navigation is verified by the experiment outlined in Section 4.2.3.
3.6 Initial Turning Direction Detection
[…]
Figure 8: Graphical representation of initial turning direction detection calculation. Green: ROI to the left of the central line, Red: ROI to the right of the central line, H: Height of the image, W: Width of the image.
$O_{F}=\frac{\operatorname*{Arg\,Max}\Biggl{[}\sum_{I_{xy}=A}^{P}I(x,y)\Biggr{]}_{P=M}^{P=B}}{\operatorname*{Arg\,Max}\Biggl{[}\sum_{I_{xy}=A}^{P}I(x,y)\Biggr{]}_{P=C}^{P=N}}$ (4)
$Initial\ Turning\ Direction=\begin{cases}left,&\text{if }O_{F}>6,\\ right,&\text{if }\frac{1}{O_{F}}>6.\end{cases}$ (5)
4 Experimental Study
Three experiments were carried out to evaluate the robustness of the proposed vision-based navigation system. The first experiment will evaluate the row-following capability of the system over long distances under varying field conditions. The second experiment examines the efficacy of crop row switching manoeuvre to navigate the robot towards the next crop row to be traversed. The third experiment will analyse the effect of the crop row switching mechanism towards the overall field scale navigation system. These three experiments will demonstrate the performance of the key components of this system: row following, row switching and the ability to scale up the row following capability to achieve field scale navigation with row switching.
4.1 Experiment 1: Long Distance Navigation
This experiment was devised to examine the performance of the proposed crop row following algorithm over long distances. A circuit of 10 crop rows was selected to test the robot’s navigation using the vision based navigation system. The robot navigated autonomously through this circuit twice, during the early growth stage and the later growth stage.
4.1.1 Ground truth crop row positions
Ten crop rows marked in Figure 9 were selected to evaluate the long-distance navigation capability of the vision based navigation system in a sugar beet field. These ten crop rows had one or more physical field variations: curved crop rows ($\bullet$), weed presence ($\blacklozenge$), tramlines ($\blacktriangle$) in the image frame and discontinuities due to missing plants($\blacksquare$). This field also had an ongoing agroforestry [22] setup with 8 rows of trees planted on the right side of the field. The 10th crop row is positioned right next to one of these agroforestry tree rows. This is an additional abnormality in the field compared to the field environments on which the crop row detection AI model was trained on [28]. The row number in Figure 9 is placed at the starting traversal point of each crop row and the traversal directions encompass all four boundaries of the field. The 10 crop rows were distributed throughout the entire field with a cumulative circuit distance totalling 2.25 km.
Figure 9: Selected crop rows for long-distance navigation experiment with field variations: curved crop rows [$\bullet$], weed presence [$\blacklozenge$], tramlines [$\blacktriangle$] in image frame and discontinuities due to missing plants[$\blacksquare$]. $[1\downarrow],[\overrightarrow{2}],[3\uparrow],[\overleftarrow{4}],[5\downarrow],[\overleftarrow{6}],[7\uparrow],[\overrightarrow{8}],[9\downarrow],[10\uparrow]$
The Hexman Mark-1 robot equipped with an EMLID Reach RS+ RTK GPS was used to record the GNSS coordinated of the 10 crop rows. The horizontal kinematic precision of the EMLID Reach RS+ RTK GPS was 7 mm + 1 ppm (1 mm error per 1km baseline increment). The robot was driven at very slow speeds (<0.3 ms-1) during the early growth stage of the crop by an expert human driver using a remote controller while the ground truth crop row positions were being recorded. A third-order polynomial spline was fitted on each of the collected crop row ground truth GNSS coordinate trajectories to create a continuous representation of the crop row ground truth position as denoted in Equation 6. The number of GNSS coordinates in each trajectory is n and the spline function $S(x)$ was created using "UnivariateSpline" method in scipy python library. This cubic polynomial was subjected to a smoothing condition stated in Equation 7 with a smoothing parameter s(=2). The weight term $w_{i}$ was set to $1$ during the smoothing condition calculation. The smoothing condition avoids overfitting of the spline effectively rejecting noise in the collected GNSS coordinates. Figure 9 shows a zoomed in view of the spline fitted among the GNSS coordinates in a selected portion of the 4th crop row.
$S(x)=\sum_{i=1}^{n}a_{i}(x-x_{i})^{3}+b_{i}(x-x_{i})^{2}+c_{i}(x-x_{i})+d_{i}$ (6)
$\sum_{i=1}^{n}w_{i}\cdot(y_{i}-\text{S}(x_{i}))^{2}\leq s$ (7)
4.1.2 Navigation Evaluation
The robot navigation was tested during both early growth stage ($\leq$ 4 leaves per plant) and the late growth stage ($\geq$ 6 leaves per plant). The robot completed the entire 10 crop row circuit during both stages resulting in a total autonomous navigation distance of 4.5 km. The GNSS coordinates of the robot were recorded during each traversal of a crop row and all the points including the ground truth coordinates were converted to Universal Transverse Mercator (UTM) for analysis. The collective root mean square error for the precision of recorded GNSS trajectories amounted to 1.8 cm. The perpendicular distance from each of the recorded points during autonomous navigation to the respective ground truth spline $S_{i}(x)$ of ith crop row was calculated. This perpendicular distance is considered as the offset of the robot from the desired crop row path during autonomous crop row following (cross-track error). The average heading error during navigation was also calculated based on the difference of the instantaneous heading angles of the ground truth spline $S_{i}(x)$ of ith crop row and the autonomous navigation trajectory. The instantaneous heading angles on the autonomous navigation trajectory was calculated by obtaining the angle of the tangent line at each of the GNSS points with respect to a third-order polynomial spline (s=0.5) fitted on the autonomous navigation trajectory based on Equations 6 and 7. The spline fitted on the autonomous navigation trajectory is indicated in orange colour in Figure 10 where $\theta$ is the instantaneous heading error.
Figure 10: Visualisation of cross-track and heading error calculations.
Table 2 summarises the cross-track error and heading error during the long-distance navigation experiment. Each crop row had a different length and other physical variations as indicated in Figure 9. The overall median cross-track was recorded at 3.32 cm while the average was recorded at 3.99 cm. Both median and average heading errors were recorded at 1.24 with the maximum recorded heading error being 3.29 from row 7 during the late growth stage. Heading and cross-track errors from late growth stage navigation recorded to be slightly higher than that of early growth stage navigation. The crop row detections during the early growth stage tend to be closer to the recorded ground truth positions due to the smaller size of the plants. The crop canopy grows larger during the late growth stage and the detected crop row could have a slight offset from the actual emergence point of the plant stem (lateral meristem region). However, these errors are negligible considering the slight difference between early and late growth stage error values. The horizontal kinematic accuracy of the GNSS sensor was 7mm + 1ppm according to the specifications of EMLID Reach RS+ RTK-GNSS module [15]. The 33 mm accuracy recorded on our system is similar to the expected performance of the Reach RS+ RTK GNSS sensor with a base station located 26km away.
Table 2: Median average cross-track and heading errors during traversal of each crop row
Row Number Row Length (m) Cross-track Error (cm) Heading Error ()
Early Stage Late Stage Early Stage Late Stage
1 207.52 4.73 4.30 0.42 1.51
2 482.11 2.85 2.48 1.43 3.21
3 217.05 3.13 3.06 0.49 1.41
4 238.14 3.85 5.02 0.26 1.24
5 140.50 3.12 3.06 1.72 1.05
6 204.08 2.56 2.47 2.57 0.74
7 202.61 4.19 3.44 0.81 3.29
8 225.51 2.75 4.08 1.13 0.03
9 167.42 2.62 3.15 1.98 0.98
10 163.70 2.60 2.86 0.26 0.28
Average 224.86 3.24 3.39 1.11 1.37
4.1.3 Field Variation Analysis
The results of the autonomous navigation trials are re-evaluated based on the individual field variations available in the traversed crop rows in Table 3. The slightly higher error values were observed in the late growth stage crop row navigation compared to the early growth stage navigation trials in this analysis as well. There was a clear difference in both heading and cross-track errors for crop rows that were curved and positioned near tramlines. The system performance has not been significantly influenced by the presence of weeds and missing plants in crop rows. The initial stress testing experiments of the proposed visual servoing controller concluded that the robot could be brought back into the crop row as long as the robot stays within 20 heading deviation from the crop row [27]. The results of this long-distance navigation experiment show that the robot’s heading deviation stays under 4 during autonomously navigating a 4.5 km distance in an arable field. To this end, this experiment reinforces the promise of the proposed vision-based crop row following pipeline to be well within the tested performance margins of the employed controller. The previous experiments conducted on vision-based crop row detection [28] concluded that the proposed perception algorithm can accurately predict the crop row positions under varying field conditions. The long-distance navigation experiment presented in this paper further confirms the robustness of our crop row detection algorithm as a reliable perception method for crop row following under varying field conditions through different growth stages of the crop.
Table 3: Median average cross-track and heading errors for each physical crop row variation
Row Variation Cross-track Error (cm) Heading Error () Overall Error
Early Stage Late Stage Early Stage Late Stage Cross-track (cm) Heading ()
Curved 3.75 3.66 0.68 2.13 3.71 1.41
Weed 3.07 3.32 1.12 1.40 3.20 1.26
Tramlines 3.72 3.84 0.93 1.86 3.78 1.39
Discontinuities 3.32 3.50 0.95 1.44 3.41 1.20
4.1.4 Discussion
There were four instances where the robot completely deviated from the traversing crop row, moving towards the adjacent crop row during the 4.5km autonomous navigation trials. The human driver overtook the control of the robot in these instances and brought the robot back into the traversing crop row in all these instances as seen in Figure 11. Each of the occurrences exposed certain limitations of the hardware and software aspects of the proposed approach. The instance a in Figure 11 caused due to one of the wheels of the robot being trapped in a pothole while traversing the crop row. As seen in Figure 2, the presence of large rock fragments is a salient feature in the sugar beet field where these experiments were conducted. While potholes are a rare occurrence in arable fields, this scenario was caused by the displacement of a large rock in the inter-row space. The two instances illustrated in frame b of Figure 11 occurred during the late growth stage traversal of row 4, which is one of the most challenging crop rows out of all the 10 crop rows. Both of these occurrences were caused by having longer fragment discontinuities in the crop row due to missing plants. Albeit our crop row detection system is able to recover the missing segments of the crop row [28], it fails when there are longer segments of discontinuities in the crop row. There were no visible plants belonging to the crop row within the image frame during these two instances causing the crop row detection algorithm to falsely detect the adjacent row as the desired traversable crop row. The scenario c of Figure 11 was also due to a similar perturbation where the early growth stage crop was not growing faster in certain parts of the crop row leading the crop row detection algorithm to attempt traversing to the adjacent crop row. The results presented in Sections 4.1.2 and 4.1.3 are calculated excluding these perturbations in the autonomous navigation trajectories. However, the effect of these instances in the data towards the overall accuracy figures presented was negligible.
Figure 11: Instances where the robot left the traversing crop row failing to recover itself towards the desired path. a: Due to a pothole in the inter-row space, b, c: Due to the large discontinuities in crop row.
The oscillations in controller response during early growth stage navigation trials were observed to be relatively higher than that of the late growth stage navigation. However, the overall error during the early growth stage navigation was relatively lower compared to the late growth stage navigation trials. This observation suggests that the controller parameters could be optimized based on the growth stage for better response during autonomous navigation. The controller overshoots could be further observed by calculating the settling distance for above-average (> 4 cm) cross-track errors. The settling distances for above-average local maximums of the cross-track error curve during early-stage and late-stage crop row following trials were plotted in Figure 12. The y-axis of this plot indicates a settling distance rather than a settling time due to the constant velocity of the robot set at 0.3 ms-1. The scatter plot confirms that the majority of the above-average cross-track errors are recovered to a below-average offset within 3 m (or 10 seconds) of traversal forwards into the crop row. This also shows that the settling distance variation exhibits a peak around 9.22 cm cross-track error region with a standard deviation of 1.93 cm for cross-track errors with larger settling distance (>3 m). About 76% of the data points belonging to this sub-distribution (mean=9.22 cm, stdev=1.93 cm) originate from 3 crop rows: 1,4 and 10. While sufficient evidence indicates that the controller can quickly respond to larger cross-track errors, 76% of the data points belonging to this sub-distribution representing settling distances over 3 m (mean=9.22 cm, stdev=1.93 cm) were observed to be originating from 3 crop rows: 1,4 and 10. All these 3 crop rows are located closer to the edge of the field [^1] Row 10 is located next to an agroforestry tree row, which could be considered as an edge of the field., where the crop rows are susceptible to having longer segments of missing plants due to foraging wildlife on the crop.
Figure 12: Variation of settling distance vs. cross-track error during crop row following (Robot Speed = 0.3 ms-1)
4.2 Experiment 2: Crop Row Switching
@@
The row switching manoeuvre presented in Section 3.5 is a three-step process that involves 6 state transitions as illustrated in Figure 6. The accuracy and the performance of each transition are individually evaluated during the experiments. The Hexman Mark-1 robot was programmed to follow a crop row and automatically detect the EOR and re-entry positions. The proposed row-switching manoeuvre is automatically started based on the EOR detection trigger. The path of the robot during the row-switching manoeuvre was tracked using the onboard RTK GNSS tracker with sub-centimetre accuracy.
@@
An experiment was set up in a selected area of 10 crop rows within a real sugar beet field. The GNSS coordinates of the 10 crop rows were recorded with sub-centimetre accuracy by driving the robot through each crop row at very slow speeds by an expert human driver. These ground truth GNSS coordinates of each crop row were then used to generate a regression line which would be used as a reference to calculate errors in autonomous navigation during row switching manoeuvre of the robot. The robot was allowed to autonomously execute the row switching manoeuvre among the 10 crop rows turning in both directions (left and right turns). A total of 18 row-switching trials were conducted during this experiment as plotted in Figure 14. The GNSS coordinates were converted to Universal Transverse Mercator (UTM) coordinates for plotting and error calculations. The errors during each state transition of each row switching manoeuvre trial are illustrated in Figure 13.a where distance errors and angular errors are normalized within a scatter plot. There are 9 x-ticks in Figure 13.a between each pair of adjacent states which represents a pair of consecutive real-world crop rows where the trial took place. A box and whisker plot of the errors for each transition is denoted in Figure 13.b. This normalised representation provides a comparative illustration of the error magnitudes in each state transition. The plots corresponding to translational steps stand above the zero line while box and whisker plots of angular transitions are centered close to zero. Table 4 presents the median errors of traversal during each stage of the row-switching manoeuvre. Normalized median percentage error ($\alpha=E_{median}/E_{max,T}$) was calculated where $E_{max,T}$ is the maximum absolute error for the type $T$ (T being a distance or an angle) across the entire manoeuvre and $E_{median}$ is the median error for each transition. The vision-based transition $A\rightarrow B$ records the highest $\alpha$ value representing the transition with the highest error. The remaining transitions record relatively lower $\alpha$ values representing accurate navigation relative to the vision-based transition.
@@
Table 4: Linear and angular errors during each transition of the row switching maneuver.
@@
4.2.1 Crop row exit and headland entry
The row exit step ($A\rightarrow B\rightarrow C$ ) in the proposed row switching manoeuvre could be identified with the densely distributed GNSS coordinates in Figure 14 leading towards headland from within the crop row in each trial. The dense distribution of GNSS coordinates attributed to the slower speed of row switching manoeuvre relative to the in-row navigation speed of the crop row navigation algorithm.
The row crop row exit ($A\rightarrow B$ transition) is a vision-based navigation stage where the headland entry ($B\rightarrow C$ transition) uses the wheel odometry to guide the robot into the headland area. This difference in feedback modalities is reflected in the distance errors of each transition. The visual feedback in $A\rightarrow B$ transition ensures that the robot has successfully reached the EOR position with visual confirmation. Therefore, it is vital for successfully reaching the EOR despite the higher error margins compared to wheel odometry. The majority of the errors in $A\rightarrow B$ and $B\rightarrow C$ transitions are positive, which indicates that the robot always travels further into the headland area beyond desired positions at states $B$ and $C$. This trend does not have significant adverse impact on the overall row-switching manoeuvre since such extra distance traversed into the headland wouldn’t cause the robot to damage crops during the U-turn step.
The robot would stop at 52.6 cm ($L_{robot}$) away from the actual EOR position in an ideal $C$ state. However, the overall maximum error in the row exit step ($E_{ABC,max}$) was recorded as 64.27 cm, a distance at which the robot would move further away from the desired position at state C. Based on these observations, the minimum width $W_{H,min}$ for the headland space was calculated to be 143.17 cm using Equation 8. The coefficient of $L_{robot}$ term in Equation 8 was set to $1.85$ since the RTK-GNSS receiver used to measure the robot motion was mounted at 45 cm behind the front of the robot. This coefficient of $L_{robot}$ term must be changed to $(1+R)$ where $R$ depends on the position of the RTK-GNSS receiver on the robot if this experiment is being repeated on a different robot. The robot was expected to be aligned with the crop row at state $A$ within a heading error margin of 2. A heading error beyond 2 at state $A$ leads the robot to cross into the headland buffer of an adjacent crop row at state $C$, which would cause it to skip 1 crop row during switching or re-enter the same crop row it traversed as seen in the unsuccessful attempts in Figure 14.